The formula 0.85x + 0.50(a-x) = 0.70a expresses the number of gallons x of a fungicide that is85% detergent and will be mixed with a 50% detergent solution to achieve a 70% detergent solution. The 70% solution will fill a 10-gallon tank. What is the value of "a" in the equation?
A. 0.7
B. 0.35
C. 1.40
D. 10

Answer by ptaylor(2198)

Now we know that the amount of pure detergent in the 85% solution (0.85x) plus the amount of pure detergent in the 50% solution (0.50(10-x)) has to equal the amount of pure detergent in the final mixture (0.70)*10, so our eq to solve is:
0.85x+0.50(10-x)=0.70*10, so a=10
Hope this helps---ptaylor